OpenStudy (anonymous):

Find the slope and the y-intercept of the line. y= -5

jimthompson5910 (jim_thompson5910):

y = -5 is the same as y = 0x + (-5)

jimthompson5910 (jim_thompson5910):

compare y = mx+b to y = 0x + (-5)

Nnesha (nnesha):

slope intercept form equation is y =mx + b where m is slope with is always with x and b is y-intercept so is there any x??

OpenStudy (anonymous):

the answer is -5

OpenStudy (anonymous):

no x intercept

OpenStudy (anonymous):

what is the slope

OpenStudy (anonymous):

it is -5

Nnesha (nnesha):

slope which is with x like m in this equation y =mx +b

Nnesha (nnesha):

so there is no X mean NO SLOPE

OpenStudy (anonymous):

ok

Nnesha (nnesha):

y = b ---> b=y-intercept no x no slope b so there is only y---intercept :